### v1

# Pressure, Reference Points, and Risk-Taking Behavior: Evidence from Bench Press Competitions

Abstract
This study explores the impact of pressure and reference points on risk-taking behavior and success outcomes in competitive settings. Using a unique panel dataset from official bench press competitions, we decompose the effects of pressure from rivals on lifters’ risk-taking decisions and realized outcomes. Our findings reveal that pressure from lower- and higher-ranked rivals drives more aggressive weight attempts, particularly in earlier stages, with the likelihood of success increasing as pressure intensifies. In third and final attempts, pressure from higher-ranked competitors leads to more conservative choices. Counterfactual analysis shows that in the absence of pressure, lifters adopt more conservative strategies in second attempts and take greater risks in third attempts, though with reduced success probabilities. These results underscore the importance of strategic focus on personal goals and the potential benefits of disregarding external pressures in the later stages of competition.
Keywords: reference point, pressure, risk, benchpress
JEL code: D90, L83, Z22, L23
1 Introduction
Understanding the behavioral effects of reference points, particularly in the context of pressure and risk-taking, is a critical area of research. The core concept is that an individual’s assessment of an outcome is influenced not only by the outcome itself but also by how it compares to a reference point (Tversky and Kahneman 1992). A significant challenge in this field is disentangling the effects of pressure from a reference point on risk-taking behavior and realized outcomes. Unlike controlled experimental settings as in Schwerter (2024), where risk-taking can be modeled as a simple binary lottery, field data often presents complexities that make such decompositions difficult. In this paper, we utilize a unique sequential competition setting and extensive panel dataset from official bench press competitions, offering a rare opportunity to explore the interplay between pressure and risk-taking behavior. This dataset enables us to observe weight attempts, lifting outcomes, rankings, and the exogenous pressure exerted by rivals over time. By examining how pressure from reference points influences subsequent risk-taking decisions, we aim to enhance the understanding of behavior under uncertainty and contribute to discussions on optimizing competition design to maximize each participant’s performance.
The primary contribution of this study is the decomposition of the effects of pressure arising from reference points into distinct elements: the choice of weight attempt (analogous to a lottery choice) and the actual probability of success (analogous to a lottery gain). We also explore alternative competition designs that manipulate the availability of pressure-related information. The structure of official bench press competitions provides an ideal context to study the relationship between reference points and risk-taking behavior within the sequential game setting of weight attempts and weight lifting stages, where exogenous pressure is present. Specifically, a weight attempt under pressure from rivals can be seen as a risk-taking behavior similar to making a lottery choice. Whether the lift is successful depends on the pressure conditional on the weight attempt and the outcomes and attempts of rivals. Through this investigation, we seek to answer an analogous question in real-world competitive settings: What would be the impact on risk-taking behavior and the distribution of expected outcomes if the pressure exerted by rivals were removed?
Our results demonstrate pressure’s significant role in influencing lifters’ behavior during competition. We find that male lifters, as well as those with higher body weight and experience, are more likely to attempt higher weights in both their second and third attempts out of the three allowed. Notably, pressure from both lower and higher-ranked rivals during the second attempt leads to increased attempt weights, reflecting lifters’ tendency to challenge themselves under competitive conditions. However, in third and final attempts, the pressure from lower-ranked rivals continues to drive higher attempts, while pressure from higher-ranked rivals results in more conservative attempts, likely due to the lifter’s perceived unlikelihood of overtaking higher-ranked competitors. This highlights the influence of both competitive positioning and loss aversion on decision-making, with cautious behavior prevailing when lifters face strong upward competition in later stages.
In terms of success probability, lifters facing pressure during their second attempts exhibit a higher likelihood of success, driven by the fear of losing rank. However, by the third attempt, fatigue and the diminishing return on rank improvement cause pressure from lower-ranked rivals to have a negligible effect on success, while pressure to outperform higher-ranked rivals remains a motivating factor.
Simulations of counterfactual scenarios, where pressure factors are removed, show that lifters are more conservative in their second attempts and take on more risk in their third attempts, often attempting weights closer to or above their best. However, success probabilities decline under these conditions, especially for second attempts, as the absence of competitive pressure reduces the effort exerted by lifters. These findings suggest that while lifters can adjust their strategy in response to pressure, maintaining focus on committing to their personal goals and ignoring external pressure may be key to maximizing success, particularly in later attempts. In broader contexts such as maximizing expected gains from risky choices in similar sequential setting, individuals should commit to their original risky decisions, especially in the later stages of competition, ignoring external pressures from competitors.
1.1 Related literature
This paper contributes to three strands of the literature: pressure from reference points, risk behavior under pressure, and competition design.
First, our paper contributes to the growing body of literature on reference points in sports competition. We follow the framework of O’Donoghue and Sprenger (2018), who suggest a variety of reference points, including the status quo (e.g., prior wealth), past experience, focal outcomes, aspirations, expectations, and other possibilities such as norms and social comparisons. However, the determination of reference points is often arbitrary and context-dependent (DellaVigna 2018), with these effects being composited and mixed in actual sports settings. [^1] Baillon et al. (2020) use Bayesian hierarchical modeling to estimate the marginal posterior distributions of six reference point rules. They found that subjects most frequently used the status quo and MaxMin as reference points. Table 1 summarizes the data, reference points, classifications, and the loss-gain domain of the utility function in related studies, although the classification is not perfect.
Salient numbers, such as round finish times in marathons (Allen et al. 2017, Soetevent 2022) and predetermined par scores in golf (Pope and Schweitzer 2011), are well-known as focal reference points. Runners adjust their effort levels to meet target times, and golf players are significantly less accurate when attempting shots ahead of a predetermined reference point, labeled as “par.”  [^2] Beyond sports competition, Rees-Jones (2018) provides evidence implying that individuals have reference-dependent preferences, with zero tax due serving as the reference point. As experimental evidence, Abeler et al. (2011) conducted an experiment manipulating the fixed pay amount as a reference point in a real-effort task. In this experiment, either fixed pay or piecework pay for a correct answer was randomly determined after task completion. They found that subjects were more likely to stop when the expected piecework pay equaled the fixed pay, exerting more effort when the fixed pay was high. Corgnet et al. (2015) also show that goal setting by managers is most effective for team workers when monetary incentives are strong. Goals set by athletes based on past or current performances are treated as reference points in the form of status quo, past experience, and expectations. For example, self-reported target times in marathons (Markle et al. 2018) and personal bests in chess (Anderson and Green 2018, González-Díaz et al. 2024) serve as natural reference points, whether or not they function explicitly as goals. Theoretically, Heath et al. (1999) and Koch and Nafziger (2016) discuss the concept of goals as reference points. In our data on bench press competition, we do not observe prevalent bunching around round weights, likely because the categories in which players participate are conditional on characteristics such as age, body weight, and gender. Additionally, unlike professional sports that provide substantial rewards, bench press competitions are amateur events with no direct monetary incentives. This setting highlights lifters’ self-motivation to surpass personal bests and rank updates based on recent outcomes, independent of financial incentives. The current rank is the reference point and each lifter has more incentive to avoid a large loss from decreasing his ranking rather than to get a relatively small gain from increasing his ranking.
In terms of pressure from rivals’ outcomes, the phenomenon of “choking under pressure”—where performance declines under stress—has been extensively studied. For instance, large-sample studies reject the hypothesis that first-mover advantage in soccer or being slightly behind increases the likelihood of winning in basketball, football, or rugby (Kocher et al. 2012, Teeselink et al. 2023). [^3] Similar studies include the effects of spectators on penalty shootouts in soccer (Dohmen 2008), free throws and playoff performance in basketball (Hickman and Metz 2015, Böheim et al. 2019, Cao et al. 2011, Toma 2017, Morgulev and Galily 2018), the impact of bookmakers’ odds on soccer play (Bartling et al. 2015), putting performance in golf (Hickman and Metz 2015), the influence of decisive moments on dart throws (Teeselink et al. 2020), and biathlon performance (Harb-Wu and Krumer 2019, Lindner 2017), the effect of previous game outcomes on current outcomes in hockey (Kniffin and Mihalek 2014), and gender differences in pressure in tennis (Cohen-Zada et al. 2017, Paserman 2023). While some papers implicitly link the cause of pressure to reference points, we do not explicitly classify pressure under the loss-gain utility framework, as the specific cases do not fit neatly within that domain. Our study also relates to peer effects in terms of externalities from rivals’ abilities and performance. Guryan et al. (2009), using random assignment in high-stakes golf tournaments, find no evidence of peer effects on average. In contrast, Yamane and Hayashi (2015) observe positive peer effects in swimming competitions, particularly suggesting that being chased by peers with lower personal bests enhances performance. These mixed results indicate that peer effects vary across different contexts, with Nishihata (2022) showing that peer effects among speed skaters depend on race distance and prize money. In our study, we leverage the sequential timing of observing rivals’ outcomes, applying the attempt weight, and executing the lift to decompose the effects of rivals’ outcomes on both the weight selection and the success probability of the lift.
Second, this paper contributes to the literature on the effect of reference points on risk-taking actions. The most closely related study is Puente-Moncayo (2019), which examines the effect of current scores on the tendency to take high-risk shots in tennis. The author follows Ely et al. (2017), using the approach developed by Klaassen and Magnus (2009) (a structural model) to show that first serves should be riskier than second serves, and using two first serves (or two second serves) is suboptimal. In the experimental economics literature, Schwerter (2024) conducted experiments on risk-taking behavior when subjects were exposed to exogenously predetermined peer earnings. The author found that subjects exposed to higher peer earnings increased their risk-taking behavior. Similarly, Post et al. (2008) highlight the relevance of risk-taking behavior in path-dependent contexts, using data from the TV show “Deal or No Deal” and a replicated experiment where subjects choose to either “deal” for a sure prize or “no deal” to continue for an uncertain prize. Our paper is more ideal than these examples for several reasons. First, the weight selection conditional on personal bests and rivals’ outcomes is analogous to choosing a lottery, where the probability of successfully lifting the weight can be inferred from data by both lifters and econometricians. Second, lifting weights involves a simple, monoarticular action, which leads to relatively fewer external factors influencing the outcomes, compared to other sports.
Third, our paper contributes to the literature on competition design aimed at maximizing each player’s score within a group. This contribution parallels general frameworks in competition design within organizations and is related to contest theory, although sabotage and uncooperative behavior, which can occur in real-world workplace competition, are less likely in sports competition. [^4] Drago and Garvey (1998) find that strong promotion incentives reduce helping behavior, even though individual effort increases. Bandiera et al. (2005) show that fruit pickers reduce their effort under relative incentives when their peers are friends, due to the negative externality of individual effort on others’ payoffs. As outlined by Lazear and Oyer (2013), relative performance evaluation (RPE) exists in two forms. The first allows firms to filter out common shocks by comparing an individual’s performance to that of a peer group. Gibbons and Murphy (1990) provide empirical evidence suggesting that RPE was used in the compensation contracts of U.S. CEOs from 1974 to 1986. The second form of RPE involves fixed rewards based on the performance rankings of participants, commonly referred to as a rank-order tournament. In tournament theory, as modeled by Lazear and Rosen (1981), large differences in payoffs across ranks motivate those at lower levels to exert more effort, while some risk-averse agents may exert less effort when the importance of winning is diminished. Rank-order tournaments have a wide range of applications, including sports competitions. For example, consistent with tournament theory, Brown (2011) shows that the presence of a superstar highly likely to win reduces the performance of other players. In our context, the unique structure of official bench press competitions allows us to compare alternative competition settings to the current format through simulations based on our estimates.
Table 1: Empirical sports and experimental papers studying reference points
Paper Data Reference Point Classification Loss-Gain Domain
Pope and Schweitzer (2011) Golf Par For Each Hole Focal Outcomes Score For Each Hole
Pope and Simonsohn (2011) Basketball Round Numbers In Performance Focal Outcomes Current Outcome
Allen et al. (2017) Marathon Round Finish Time Focal Outcomes Finish Time
Essl and Jaussi (2017) Lab Experiment Points (ECU) When <,><,>< , > 10sec Focal Outcomes Points From Real-Effort Task
Soetevent (2022) Running Event Round Finishing Time Focal Outcomes Finishing Time
Schwerter (2024) Lab Experiment Peer Earnings Focal Outcomes Earnings From Binary Lotteries
Post et al. (2008) TV Shows Earlier Expectations Expectations Bank Offer
Abeler et al. (2011) Lab Experiment Amount Of Fixed Pay Expectations Piece Rate For Correct Answers
Bartling et al. (2015) Soccer Expected Outcome By Odds Expectations Current Outcome
Markle et al. (2018) Marathon Self-Reported Target Time Status Quo Finish Time
Kniffin and Mihalek (2014) Hockey The Outcomes For Game 1 Past Experience Score In Game 2
Anderson and Green (2018) Chess Online Personal Best Rating Past Experience Rating
González-Díaz et al. (2024) Chess In-Person Personal Best Rating Past Experience Rating
Berger and Pope (2011) Basketball Performance Of The Opponent Social Comparison Current Outcome
Puente-Moncayo (2019) Tennis Rival’s Score Social Comparison Score
Teeselink et al. (2023) Four Sports Performance Of The Opponent Social Comparison Current Outcome
This Paper Bench Press Current Own Ranking Social Comparison Ranking
Baillon et al. (2020) Lab Experiment 6 Rules Mixed Lottery Options
Notes: We follow the referent list of O’Donoghue and Sprenger (2018) which suggest a number of possibilities: the status quo (e.g., prior wealth), past experience, focal outcomes, aspirations, expectations, and other candidates including norms and social comparisons, although the reference point determination is, in ways, arbitrary and context-dependent (DellaVigna 2018) and these effects are composited and mixed in the actual sports.
6 rules (Status Quo, MaxMin, MinMax, X at Max P, Expected Value, Prospect Itself), ECU=Experimental Currency Unit.
2 Data
2.1 Data source
We use the data from OpenPowerlifting which is a community service project to create a permanent and open archive of the world’s powerlifting data. [^5] https://gitlab.com/openpowerlifting. The OpenPowerlifting database includes a wide range of data fields of official powerlifting competition. It includes competition details such as federation name, date and location, level (local, national, international), athlete information such as name, age, weight class, gender, lift results of squat, bench press, and deadlift (with attempted weights and success/failure status), Wilks score and other relative strength metrics, and rankings of placement within the competition historical rankings and records.
In this paper, we focus on the bench press competition which is the most popular division and less complex than Squat-Bench-Deadlift (SBD) composite competition. We use all official attempts data of 15-69 aged lifters at competition categories based on age-class, weight-class, and gender, in which the number of participant lifters is at least two to capture competition pressure.
Table 2: Summary Statistics
Equipment N mean sd min max
Raw Male 175829 0.86 0.35 0.00 1.00
Personal best 175829 74.73 81.06 0.00 307.50
First attempt weight 175829 130.96 44.77 5.00 330.00
Second attempt weight 175829 137.42 45.89 20.00 352.50
Third attempt weight 175829 141.62 46.56 20.00 352.50
Successful first attempt 175829 0.86 0.35 0.00 1.00
Successful second attempt 175829 0.72 0.45 0.00 1.00
Successful third attempt 175829 0.40 0.49 0.00 1.00
Best attempt 168959 137.97 46.24 0.00 325.00
Age 175829 31.77 11.21 13.00 69.00
Bodyweight 175829 85.35 20.37 24.90 240.00
Single-ply Male 76605 0.80 0.40 0.00 1.00
Personal best 76605 75.12 92.05 0.00 402.50
First attempt weight 76605 156.17 64.67 5.00 445.00
Second attempt weight 76605 162.84 65.51 20.00 445.00
Third attempt weight 76605 168.06 66.65 20.00 445.00
Successful first attempt 76605 0.77 0.42 0.00 1.00
Successful second attempt 76605 0.65 0.48 0.00 1.00
Successful third attempt 76605 0.40 0.49 0.00 1.00
Best attempt 70687 160.23 64.68 0.00 445.00
Age 76605 32.32 12.93 13.00 69.00
Bodyweight 76605 84.90 23.19 30.50 245.00
Multi-ply Male 14628 0.97 0.18 0.00 1.00
Personal best 14628 79.40 107.58 0.00 425.00
First attempt weight 14628 197.02 63.56 25.00 480.00
Second attempt weight 14628 205.79 64.59 30.00 500.50
Third attempt weight 14628 212.03 65.67 30.00 520.00
Successful first attempt 14628 0.70 0.46 0.00 1.00
Successful second attempt 14628 0.57 0.49 0.00 1.00
Successful third attempt 14628 0.35 0.48 0.00 1.00
Best attempt 12820 201.35 63.09 0.00 430.00
Age 14628 32.17 10.08 13.00 69.00
Bodyweight 14628 94.65 19.64 34.40 207.80
Sources: The OpenPowerlifting database. We also observe Wilks score and other relative strength metrics, and rankings of placement within the competition historical rankings and records used for robustness check.
Table 2 presents summary statistics across three types of powerlifting equipment: raw, single-ply, and multi-ply. Each category includes attributes such as personal best, first, second, and third attempt weights, successful attempts, best attempt, age, and body weight. The raw category has the largest number of participants, with 175,829 individuals, an average personal best of 74.73 kg, and a mean bodyweight of 85.35 kg. The single-ply category includes 76,605 participants, with an average personal best of 75.12 kg and a mean bodyweight of 84.90 kg. In the multi-ply category, there are 14,628 participants, with an average personal best of 79.40 kg and a mean bodyweight of 94.65 kg.
Across all equipment types, the mean weights lifted increase from the first to the third attempts, and the percentage of successful attempts decreases with each successive attempt. This indicates that while lifters attempt heavier weights in successive attempts, the success rate declines. Approximately 80% of participants are male. The average age of participants is consistently around the early 30s across all categories. These statistics highlight key performance trends and demographic characteristics in competitive powerlifting, emphasizing differences in performance and success rates among different equipment types and attempts.
Personal best weights of lifters who have never participated in an official bench press competition are recorded as zero. In the empirical analysis, we focus on the second and third attempts and treat the realized outcome in the first attempt as the personal best for such lifters because the success probability in the first attempt is more than 90%.
2.2 Game setting
The objective of each lifter is to lift the maximum weight possible from a prone position on a flat bench to improve the own rank. In case of a tie (same weight lifted), the lighter lifter or the lifter who achieved the weight first (depending on the federation’s rules) is ranked higher. Competitors are divided into categories based on weight class, age class, equipment class, and gender to ensure fair competition.
Competition Structure
Competitors declare their opening weight (first attempt) before the competition starts. The declared weights determine the initial lifting order. Lifters are arranged in ascending order based on their declared opening attempts. That is, the lifter with the lightest weight goes first, followed by the next lightest, and so on.
Each lifter has three attempts.666 Lifting Procedure for each lifter at each attempt is divided into the following. 1. Setup: The lifter lies on their back on the bench, with feet flat on the floor or on the bench’s footrests, ensuring contact with the bench throughout the lift. 2. Grip: The lifter grips the barbell, usually slightly wider than shoulder-width apart. 3. Unracking: With the help of spotters if needed, the lifter unracks the barbell and holds it with arms fully extended above their chest. 4. Lowering: The barbell is lowered in a controlled manner to touch the chest, ensuring it is stable and paused for a moment. 5. Press Command: The judge gives a “Press” command once the barbell is motionless on the chest. The lifter then presses the barbell back to the starting position with arms fully extended. 6. Rack Command: Upon completion, and with the barbell in a stable position, the judge gives the “Rack” command, signaling the lifter to return the barbell to the rack. Each attempt is judged based on control, stability, and completion of lifting. A lift is considered successful if the lifter adheres to all the rules and completes the lift as per the judge’s commands from the viewpoints of two of three referees. After completing an attempt, the lifter must declare the weight for their next attempt within a specific time frame (usually 1 minute after their attempt). Once a weight is declared, it generally cannot be decreased, only increased. The order of lifters is adjusted after each round of attempts based on the declared weights for the next attempt, again proceeding from lightest to heaviest. This ensures that the competition is efficient and fair.
Concretely, in the first attempt, all lifters complete their first attempt in the sequential order of the lightest to heaviest declared weights. In the second attempt, the order is recalculated based on the declared weights for the second attempt. Lifters who declared lighter weights for their second attempt will lift first. In the third attempt, the order is again recalculated for the third attempt based on the new declared weights. If a lifter fails an attempt, they can choose to reattempt the same weight or increase it for their next attempt. The order will still be based on the declared weight for the next attempt, regardless of whether the previous attempt was successful.
Each lifter’s highest successful attempt is recorded as their final score. If two lifters achieve the same highest lift, the lighter lifter or the lifter who achieved the weight first (depending on the federation’s rules) ranks higher.
Example of Attempt Order:
Figure 1: The structure of the first and second attempts: Three lifters case
Notes: The box means that the player appears on the stage. $W_{it}^{j}$ is lifter $i$’s attempt weight in $j$-th attempt in competition $t$.
To highlight the available information of each lifter at each attempt, consider a competition with three lifters (A, B, and C). The timeline of the first and second attempts is shown in Figure 1. For opening attempts, lifters $A,B$, and $C$ declare 100 kg, 105 kg, and 110 kg before the competition. Then, the order will be $A,B,C$. Lifter $A$ attempts $100$kg then the outcome is realized. Next, lifter $B$ attempts $105$kg. At the same time, lifter $A$ declares the next attempt weight $110$kg in one minute conditional on $A$’s own first outcome and $B$ and $C$’s attempt weights before $B$’s outcome is realized because lifter $B$ is trying 105kg at that time. Lifter $B$’s outcome is realized. Next, lifter $C$ attempts $110$kg. At the same time, lifter $B$ declares the next attempt weight $115$kg in one minute knowing $B$’s own first outcome, $A$’s outcome and next attempt, and $C$’s attempt weight before $C$’s outcome is realized because lifter $C$ is trying 110kg at that time. Lifter $C$’s outcome is realized. Lifter $C$ declares the next attempt weight $112.5$kg in one minute knowing $A$ and $B$s’ outcomes, their next attempts, and $C$’s own first outcome. For the second attempts, the new order will be A (110 kg), C (112.5 kg), B (115 kg). After the same procedure, the new order for the third attempts, will be similarly determined by the declared weights at the second attempts.
2.3 Exogenous pressure
The above sequential game structure in which each player does not know the subsequent attempts and outcomes provides exogenous variations to the choice of the attempt weight and the success of lifting the attempt weight, which is analogous to the lottery choice and realized gain of the lottery. We focus on the moves of the closest rivals to each lifter, that is, his lower rank and higher rank rival lifters, which are critical to finalize his rank in the competition.
The choice of attempt weight
In the choice of attempt weight at the second attempt, first, each lifter knows the first attempt outcome and the second attempt weight of his one lower rank rival, which are uncorrelated with his unobserved characteristics such as his body condition at the competition date. The exogenous pressure affects the risk of being turned around by one lower rank rival, that is, the risk of loss. Second, each lifter also knows the first attempt outcome of his one higher rank rival and expects the second attempt of the rival, which is uncorrelated with his unobserved characteristics. We assume that each lifter has a rational expectation of the higher rank rival’s attempt because it is predictable so that we treat the second actual attempt of the rival as the second actual attempt known to the lifter. As the first attempt weight of his one higher rank rival heavier than his current best weight, it is more difficult to turn around the rival. The exogenous pressure affects the risk of challenging turning around his one higher rank rival, that is, the risk of gain.
The success of lifting the attempt weight
In the first attempt of lifting the attempt weight, first, each lifter knows the first attempt outcome and the second attempt weight of his one lower rank rival, which are uncorrelated with his unobserved characteristics such as his body condition at the competition date. The exogenous pressure affects the risk of being turned around by one lower rank rival, that is, the risk of loss. If his second attempt fails, he is turned around the rival. Second, each lifter also knows the second attempt weight of his one higher rank rival, which is also uncorrelated with his unobserved characteristics. If his second attempt is successful, he has more chances to turn around the rival. The exogenous pressure affects the risk of challenging turning around his one higher rank rival, that is, the risk of gain.
2.4 Data pattern
We illustrate data patterns of the choice of attempt weight and the success probability of the attempt weight conditional on the above exogenous pressure. For illustration, we show the case of the most popular competition category (Male, Raw, 24-39 age class). The data pattern is prevalent in the other competition categories, shown in the Appendix. We omit the data pattern for the first bench press attempts because the success probability at the stage is more than 75%. [^7] See Appendix for all plots for all competition categories for reference.
Attempt weight, success probability, and pressure
Figure 2: Success rate conditional on pressures at second attempt (Male, Raw, 24-39 age class)
Notes: The distance from the current best was rounded to the nearest integer. Distance categories with five or fewer lifters were excluded from the plot to avoid skewing the success probabilities to 0 or 1.
Figure 2 displays success probabilities for second bench press attempts across different weight classes, showing the highest success rates when the attempted weight is slightly below the lifter’s personal best. Across almost all classes, success peaks around a distance of zero, with success probabilities typically ranging from 0.75 to 1.00 at this point. As the attempted weight deviates from the best, success rates drop significantly, often falling below 0.25 at more extreme deviations, particularly for heavier attempts. In heavier weight classes (120 kg and 120+ kg), the range of successful attempts is slightly broader, but the general trend of declining success with larger deviations remains consistent across all weight categories. While both pressured (“turning around” or “turned around”) or non-pressured groups follow similar patterns, the pressured group generally maintains slightly higher success probabilities, especially near the lifters’ personal best.
Figure 3: Success rate conditional on pressures at third attempt (Male, Raw, 24-39 age class)
Notes: The distance from the current best was rounded to the nearest integer. Distance categories with five or fewer lifters were excluded from the plot to avoid skewing the success probabilities to 0 or 1.
Figure 3 illustrates more remarkable patterns for third bench press attempts. In comparison to the second attempt, the success rates in the third attempt are noticeably lower across all weight classes. While success rates during the second attempt peaked between 0.75 and 1.00 near the lifters’ personal best, the third attempt generally shows lower peak success rates, ranging between 0.5 and 0.8. Additionally, the drop in success rates as lifters deviate from their personal best is more pronounced in the third attempt, with success probabilities frequently falling below 0.25 at larger deviations. The distinction between the “turning around” groups (0 and 1) remains consistent, though the overall lower success rates make the difference less noticeable in the third attempt. Overall, lifters appear to struggle more in the third attempt, especially with larger deviations from their best, indicating a significant drop in success probability compared to the second attempt.
In summary, the above data patterns imply that when players are under pressure in some critical situation to change their own rank, they attempt more challenging weights and achieve them with higher probability.
Who are the rivals?
We examine the pressure exerted by rivals during competition. A natural question arises: who constitutes a rival for each lifter within a competition? Table 3 illustrates the distribution of rank changes for lifters between their second and third attempts. The majority of lifters maintain their rank, as shown by the high proportion of zero rank changes (68% in the second attempt and 82% in the third attempt). A smaller percentage of lifters experience a rank increase ($+1$) between attempts (15% in the second and 10% in the third), while the proportion of those gaining two or more ranks is notably lower (6% in the second and 2% in the third). Similarly, lifters dropping one rank ($-1$) account for 7% in the second attempt and 5% in the third, with even fewer dropping two or more ranks (4% and 2%, respectively). Based on this distribution, we focus on the immediate competitors, specifically the lifter ranked directly below and the one ranked directly above, as these rivals most significantly influence rank changes.
Table 3: Distribution of rank changes
Changes in ranking 2nd 3rd
$\leq-2$ 0.04 0.02
$-1$ 0.07 0.05
00 0.68 0.82
$+1$ 0.15 0.10
$+2\leq$ 0.06 0.02
Notes: We compute the distribution of rank changes using all achieved outcomes of all lifters in our data.
3 Theoretical prediction
We provide theoretical predictions based on expected utility and prospect theory. [^8] See the survey of O’Donoghue and Sprenger (2018) and DellaVigna (2018) for reference. The choice of attempt weight is analogous to a binary lottery choice from a set of lotteries ${(x,q(x))}$, where $x$ is the attempt weight and $q(x)\in[0,1]$ is the objective probability of success for the attempt, which is naturally assumed to be a decreasing function in $x$. The expected evaluation of lottery $(x,q(x))$ is
$U(x,q(x)|r)=v(x-r)q(x),$
where $v(\cdot)$ captures reference-dependent utility of outcome $x$ given reference point $r$. Following Tversky and Kahneman (1992), $v(\cdot)$ is specified as
$v(x-r)=\left\{\begin{array}[]{ll}(x-r)^{\alpha}&\text{if }x\geq r\\ -\lambda(r-x)^{\alpha}&\text{if }x<r.\end{array}\right.$
where $0<\alpha\leq 1$ and $\lambda>1$ capture diminishing sensitivity—concave curvature in the gain domain and a convex curvature in the loss domain—and loss aversion, respectively.
3.1 Lifting stage
Consider the attempt and lifting stages backward. We do not incorporate strategic decisions across stages to avoid complications making the situation intractable, but this would be an interesting research avenue. In the lifting stage, the reference point $r$ is the lifter’s current ranking, normalized to $r=0$, and a change in ranking induces $x\in\{-1,0,1\}$, so $\alpha$ does not matter in this setting. Let $p_{H}$ denote the ex-ante probability of the higher rival’s success, which is unknown to lifter $i$ in the lifting stage. Also, let $p_{L}$ denote the ex-ante probability of the lower rival’s success, which is unknown to lifter $i$ in the attempt stage.
Table 4 describes the utility tuple $(v_{1}(x),(v_{2}(x),v_{3}(x)))$ conditional on four scenarios determined by pressure from lower and higher rivals, where $v_{1}(x)$ is the utility when lifter $i$’s attempt weight $x$ fails, and $(v_{2}(x),v_{3}(x))$ is the utility when lifter $i$’s attempt is successful and the higher rival’s attempt either succeeds or fails. If there is no pressure from either lower or higher rivals (i.e., (No, No)), the current lifting attempt does not change lifter $i$’s rank whether successful or not, meaning $(v_{1}(x),(v_{2}(x),v_{3}(x)))=(0,(0,0))$. If there is pressure only from a lower rival (i.e., (Yes, No)), lifter $i$ experiences disutility $-\lambda(<-1)$ upon failure, but zero otherwise. If there is pressure only from a higher rival (i.e., (No, Yes)), lifter $i$ gains 1 if the higher rival fails, with the probability of $1-p_{H}$. If there is pressure from both lower and higher rivals, lifter $i$ experiences disutility $-\lambda(<-1)$ upon failure, gains 00 when both lifter $i$ and the higher rival succeed, and gains 1 if the higher rival fails, with the probability of $1-p_{H}$. Based on these scenarios, lifter $i$ has the strongest incentive to make an effort to avoid loss and gain when there is pressure from his lower and higher rivals.
Table 4: Utility tuple $(v_{1}(x),(v_{2}(x),v_{3}(x)))$
Lower rival’s outcome turning around $i$ Higher rival’s attempt is turned around by $i$ No Yes No $(0,(0,0))$ $(0,(0,1))$ Yes $(-\lambda,(0,0))$ $(-\lambda,(0,1))$
3.2 Attempt stage
Next, consider the attempt stage. Let $x_{L}$ denote the lower rival’s actual attempt weight which is observable to lifter $i$ during the attempt stage. Also, let $x_{H}$ denote the higher rival’s expected attempt weight under lifter $i$’s rational expectations. Although this is not observable to lifter $i$’s attempt stage, we treat this as observable or highly predictable information to lifter $i$ [^9] Using our data, we discuss the predictability of higher rival’s attempt weight given observable information at lifter $i$’s attempt stage in Section 7. Additionally, let $y_{L}$, $y_{H}$, and $y_{i}$ represent the best outcome weights from earlier attempts within the competition for the lower rival, the higher rival, and lifter $i$, respectively, which are known in lifter $i$’s current attempt stage. The inequality conditions $y_{L}<y_{i}<y_{H}$, determined by rank order, and $y_{L}<x_{L}$, $y_{i}<x_{i}$, and $y_{H}<x_{H}$, meaning that the outcome weight must not exceed the corresponding attempt weight, must hold. Although we can assume that $x_{L}<x_{H}$, meaning that the lower rival does not attempt to overtake two higher-ranked rivals as inferred from our data, must hold, we keep generality and allow $x_{L}>x_{H}$. These conditions determine the joint probability of the outcome tuple as shown in Table 5.
Table 5: Joint probability of outcome tuple
Pressure $i$ $L$ $H$ Outcome $i$’ rank Probability
$x_{i}<\min\{x_{L},x_{H}\}$ S S S $x_{i}<\min\{x_{L},x_{H}\}$- $q(x)p_{L}p_{H}$
$x_{i}<\min\{x_{L},x_{H}\}$ S S F $x_{i}<\min\{x_{L},y_{H}\}$- $q(x)p_{L}(1-p_{H})P(x_{i}<y_{H})$
$x_{i}<\min\{x_{L},x_{H}\}$ S S F $y_{H}<x_{i}<x_{L}$ K $q(x)p_{L}(1-p_{H})P(x_{i}>y_{H})$
$x_{i}<\min\{x_{L},x_{H}\}$ S F S $y_{L}(<y_{i})<x_{i}<x_{H}$ K $q(x)(1-p_{L})p_{H}$
$x_{i}<\min\{x_{L},x_{H}\}$ S F F $y_{L}(<y_{i})<x_{i}<y_{H}$ K $q(x)(1-p_{L})(1-p_{H})P(x_{i}<y_{H})$
$x_{i}<\min\{x_{L},x_{H}\}$ S F F $y_{L}<y_{H}<x_{i}$ + $q(x)(1-p_{L})(1-p_{H})P(x_{i}>y_{H})$
$x_{i}<\min\{x_{L},x_{H}\}$ F S Any $y_{i}<\min\{x_{L},y_{H}\}$- $(1-q(x))p_{L}$
$x_{i}<\min\{x_{L},x_{H}\}$ F F Any $y_{L}<y_{i}<y_{H}$ K $(1-q(x))(1-p_{L})$
$x_{L}<x_{i}<x_{H}$ S S S $x_{L}<x_{i}<x_{H}$ K $q(x)p_{L}p_{H}$
$x_{L}<x_{i}<x_{H}$ S S F $x_{L}<x_{i}<y_{H}(<x_{H})$ K $q(x)p_{L}(1-p_{H})P(x_{i}<y_{H})$
$x_{L}<x_{i}<x_{H}$ S S F $\max\{x_{L},y_{H}\}<x_{i}$ + $q(x)p_{L}(1-p_{H})P(x_{i}>y_{H})$
$x_{L}<x_{i}<x_{H}$ S F S $y_{L}(<x_{L})<x_{i}<x_{H}$ K $q(x)(1-p_{L})p_{H}$
$x_{L}<x_{i}<x_{H}$ S F F $y_{L}<x_{i}<y_{H}(<x_{H})$ K $q(x)(1-p_{L})(1-p_{H})P(x_{i}<y_{H})$
$x_{L}<x_{i}<x_{H}$ S F F $y_{L}<y_{H}<x_{i}$ + $q(x)(1-p_{L})(1-p_{H})P(x_{i}>y_{H})$
$x_{L}<x_{i}<x_{H}$ F S S $y_{i}<x_{L}<x_{H}$- $(1-q(x))p_{L}p_{H}P(x_{L}>y_{i})$
$x_{L}<x_{i}<x_{H}$ F S S $x_{L}<y_{i}<x_{H}$ K $(1-q(x))p_{L}p_{H}P(x_{L}<y_{i})$
$x_{L}<x_{i}<x_{H}$ F S F $x_{L}<y_{i}<y_{H}$ K $(1-q(x))p_{L}(1-p_{H})P(x_{L}<y_{i})$
$x_{L}<x_{i}<x_{H}$ F S F $y_{i}<\min\{x_{L},y_{H}\}$- $(1-q(x))p_{L}(1-p_{H})P(x_{L}>y_{i})$
$x_{L}<x_{i}<x_{H}$ F F S $y_{L}<y_{i}<x_{H}$ K $(1-q(x))(1-p_{L})p_{H}$
$x_{L}<x_{i}<x_{H}$ F F F $y_{L}<y_{i}<y_{H}$ K $(1-q(x))(1-p_{L})(1-p_{H})$
$\max\{x_{L},x_{H}\}<x_{i}$ S Any Any $\max\{x_{L},x_{H}\}<x_{i}$ + $q(x)$
$\max\{x_{L},x_{H}\}<x_{i}$ F S Any $x_{L}<y_{i}<\min\{y_{H},x_{H}\}$ K $(1-q(x))p_{L}P(x_{L}<y_{i})$
$\max\{x_{L},x_{H}\}<x_{i}$ F S Any $y_{i}<\min\{x_{L},y_{H}\}$- $(1-q(x))p_{L}P(x_{L}>y_{i})$
$\max\{x_{L},x_{H}\}<x_{i}$ F F Any $y_{L}<y_{i}<\min\{y_{H},x_{H}\}$ K $(1-q(x))(1-p_{L})$
Notes: F = failed, S = success, K = keep. We ignore the equality case like $x_{L}=y_{i}$ in the probability because the body weight information works as a tie-breaker for the case.
Given the values of $x_{L},x_{H},y_{L},y_{H}$, and $y_{i}$ with utility tuple in Table 4, the expected evaluation of the lottery $(x_{i},q(x_{i}))$ for lifter $i$ with reference point $r_{i}$ is
$U(x_{i},q(x_{i})|r_{i})=v(x_{i}-r_{i})q(x_{i})$
=$=\left\{\begin{array}[]{ll}\Bigl{(}q(x_{i})p_{L}p_{H}+q(x_{i})p_{L}(1-p_{H})P(x_{i}<y_{H})+(1-q(x))p_{L}\Bigr{)}[-\lambda]&\text{if }x_{i}<x_{L},x_{H}\\ +\Bigl{(}q(x_{i})(1-p_{L})(1-p_{H})P(x_{i}>y_{H})\Bigr{)}[1]\\ \Bigl{(}(1-q(x_{i}))p_{L}p_{H}P(x_{L}>y_{i})+(1-q(x_{i}))p_{L}(1-p_{H})P(x_{L}>y_{i})\Bigr{)}[-\lambda]&\text{if }x_{L}<x_{i}<x_{H}\\ +\Bigl{(}q(x_{i})p_{L}(1-p_{H})P(x_{i}>y_{H})+q(x_{i})(1-p_{L})(1-p_{H})P(x_{i}>y_{H})\Bigr{)}[1]\\ \Bigl{(}(1-q(x_{i}))p_{L}P(x_{L}>y_{i})\Bigr{)}[-\lambda]+q(x_{i})[1]&\text{if}x_{L},x_{H}<x_{i}\end{array}\right.$ (6)
=$=\left\{\begin{array}[]{ll}\Bigl{(}q(x_{i})p_{L}p_{H}+q(x_{i})p_{L}(1-p_{H})P(x_{i}<y_{H})+(1-q(x))p_{L}\Bigr{)}[-\lambda]&\text{if }x_{i}<x_{L},x_{H}\\ +\Bigl{(}q(x_{i})(1-p_{L})(1-p_{H})P(x_{i}>y_{H})\Bigr{)}[1]\\ \Bigl{(}(1-q(x_{i}))p_{L}P(x_{L}>y_{i})\Bigr{)}[-\lambda]+\Bigl{(}q(x_{i})(1-p_{H})P(x_{i}>y_{H})\Bigr{)}[1]&\text{if }x_{L}<x_{i}<x_{H}\\ \Bigl{(}(1-q(x_{i}))p_{L}P(x_{L}>y_{i})\Bigr{)}[-\lambda]+q(x_{i})[1]&\text{if}x_{L},x_{H}<x_{i}\end{array}\right.$ (11)
Although the representation of the expected evaluation of the lottery in Equation (11) may appear complex, in particular, when $x_{L}<x_{i}<x_{H}$, it reveals that there are discontinuities at $x_{L}$ and $x_{H}$. Additionally, it is crucial to note that $P(x_{i}>y_{H})$ (that is, chance of the higher rival’s being turned around by the lifter) and $P(x_{L}>y_{i})$ (that is, chance of the lower rival’s turning around the lifter) are key factors that capture the pressure exerted by lower and higher rivals, respectively. [^10] Even without further specification, Equation (11) offers theoretical predictions regarding the optimal choice of attempt weight $x_{i}$ for each domain of $x_{i}$. First, if lifter $i$ aims to choose $x_{i}$ such that $\max\{x_{L},x_{H}\}<x_{i}$, the optimal choice for $x_{i}$ is $x_{H}+\kappa$, where $\kappa$ is the minimum weight increment as specified by the competition rules, given that $q(x_{i})$ is a decreasing function of $x_{i}$. Second, if lifter $i$ chooses $x_{i}$ such that $x_{i}<\min\{x_{L},x_{H}\}$, the optimal choice is the $x_{i}$ that maximizes $q(x_{i})P(x_{i}>y_{H})$, which leads to $x_{i}=y_{H}+\kappa$. Third, if lifter $i$ selects $x_{i}$ such that $x_{L}<x_{i}<x_{H}$, the optimal choice is $x_{i}$ that maximizes $q(x_{i})(1-p_{H})P(x_{i}>y_{H})$, which also results in $x_{i}=y_{H}+\kappa$. Further analysis to precisely determine the analytical optimal attempt weight $x_{i}$ across all domains of attempt weights would require additional specifications for $q(\cdot),p_{H},p_{L},\lambda$, and other parameters, which increases the complexity of the modeling, identification, and estimation processes, and is beyond the scope of this study and our data.
4 Estimation
4.1 Choice of attempt weight
As the first empirical exercise, we employ a linear regression model to estimate the effect of pressure on the choice of attempt weight. We regress the outcome on the observed characteristics as follows:
$\tilde{W}_{it}^{k}=X_{it}\beta+Z_{it}^{k}\gamma+\varepsilon_{it},$ (12)
where $\tilde{W}_{it}^{k}$ is the difference between lifter $i$’s attempt weight compared to his current best in competition $t$ at a $k$-th attempt, $X_{it}$ is observed characteristics of lifter $i$ and competition $t$ including gender, body weight, number of competition experiences, a dummy variable for first participation, and fixed effects for competition $t$’s equipment category, age class, division, weight class, and federation, $Z_{it}^{k}$ is a pressure variable during attempt defined later, $\varepsilon_{it}$ is assumed to be an i.i.d. error, and $\beta$ and $\gamma$ are vectors of coefficient parameters of $X_{it}$ and $Z_{it}^{k}$. Large $\tilde{W}_{it}^{k}$ implies his choosing a more challenging attempt weight. Our primary interest lies in $\gamma$, that is, the sensitivity to pressure.
The pressure variables, $Z_{it}^{k}$, which are exogenous due to the sequential game setting, represent the difference between the lifter’s current realized outcome and the attempted weight of the lower-ranked competitor at the next attempt, and the difference between the lifter’s current realized outcome and the expected attempt weight of the higher-ranked competitor. These variables capture the potential for the lifter’s rank to change by overtaking or being overtaken by the rivals.
4.2 Success probability
As the second empirical exercise, we use a linear probability model to estimate the probability of successfully lifting the attempt weight. We regress the outcome on the observed characteristics as follows:
$Y_{it}=X_{it}\beta+\tilde{Z}_{it}^{k}\gamma+\tilde{W}_{it}^{k}\delta+\eta_{it},$ (13)
where $Y_{it}$ is whether the attempt is successful ($Y_{it}=1$) or not ($Y_{it}=0$), $\tilde{Z}_{it}^{k}$ is a pressure variable during lifting, $\eta_{it}$ is assumed to be an i.i.d. error, and $\delta$ represents the coefficient for the difference between lifter $i$’s attempt weight and his current best. Note that $\tilde{W}_{it}^{k}$, the choice of attempt weight, may be correlated with unobserved body conditions in $\eta$, potentially causing endogeneity issues. To address this, we employ a two-stage least squares (2SLS) regression, using Equation (12) as the first-stage regression, with the exogenous pressure variables during the attempt, $Z_{it}^{k}$, serving as instrumental variables (IV).
The pressure variable $\tilde{Z}_{it}^{k}$ during the lift is equal to one if the attempt involves overtaking a higher-ranked rival or being overtaken by a lower-ranked rival, and zero otherwise. These variables capture the potential for rank changes through overtaking or being overtaken.
5 Results
5.1 Choice of attempt weight
Table 6: Regression of the difference between the second and third attempt weight compared to the current best on pressure
(1) (2)
Dependent Variable $W_{it}^{2}-$(best) $W_{it}^{3}-$(best)
Male -1.140 5.232
(0.424) (0.828)
Body weight 0.252 0.385
(0.025) (0.036)
Num experience 0.214 0.337
(0.102) (0.131)
1(first participation) 33.022 34.819
(1.526) (1.647)
Pressure, lower rival, 2nd 0.264
Pressure, higher rival, 2nd 0.264
Pressure, lower rival, 3rd 0.502
Pressure, higher rival, 3rd -0.092
Control X X
Num.Obs. 267062 267062
R2 0.570 0.375
R2 Adj. 0.568 0.372
RMSE 34.46 41.87
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, and federation fixed effects. The standard errors in the brackets are clustered at federation categories. $W_{it}^{2}$ and $W_{it}^{3}$ are lifter $i$’s attempt weights in the second and third attempts in competition $t$. Pressure from lower rival at 2nd is (lower rival 2nd attempt) - (his 1st outcome). Pressure from lower rival at 3rd is (lower rival 3rd attempt) - (his 2nd outcome). Pressure from higher rival at 2nd is (higher rival 2nd attempt) - (his 1st outcome). Pressure from higher rival at 3rd is (higher rival 3rd attempt) - (his 2nd outcome).
Table 6 presents the regression results for the difference between the second and third attempt weights compared to the current best, with respect to various pressure variables. Male lifters show mixed results, with a statistically significant negative coefficient for the second attempt (-1.140) and a positive coefficient for the third attempt (5.232). This suggests that male lifters are less likely to challenge in the second attempt but are more inclined to take risks in the third attempt. Body weight and experience are positively associated with higher attempt weights in both attempts, indicating that heavier and more experienced lifters tend to choose higher weights. The positive coefficients for first participation (33.022 and 34.819) indicate that first-time participants tend to choose significantly higher attempt weights compared to their current best, which is substituted by the achieved outcome in the first attempt.
For the second attempt, pressure from both lower and higher rivals shows a statistically significant positive effect (0.264 for both), suggesting that the presence of rivals encourages lifters to choose higher weights. In contrast, for the third attempt, the pressure from lower rivals has a positive association with increased weights (0.502), while pressure from higher rivals has a negative effect (-0.092). This negative coefficient implies that when the current score of a lifter is distant from that of a higher-ranked rival during the third attempt, the lifter tends to adopt a more conservative strategy by selecting a lower weight.
5.2 Success probability of lifting the attempt weight
Table 7: Regression of success probability of second and third attempts on pressure
(1) (2) (3) (4)
Dependent Variable 1(success 2nd) 1(success 3rd) 1(success 2nd) 1(success 3rd)
Male -0.009 0.006 -0.005 0.006
(0.006) (0.006) (0.005) (0.006)
Body weight -0.002 -0.002 -0.001 -0.002
(0.000) (0.000) (0.000) 1 (0.000)
Num experience 0.001 -0.001 0.001 -0.001
(0.000) (0.000) (0.000) (0.000)
1(first participation) 0.006 0.003 0.031 0.002
(0.003) (0.002) (0.004) (0.004)
$W_{it}^{2}-$(best) -0.001 -0.002
(0.000) (0.000)
$W_{it}^{3}-$(best) 0.000 0.000
(0.000) (0.000)
1(Turned around, 2nd) 0.072 0.082
(0.009) (0.009)
1(Turning around, 2nd) 0.037 0.038
(0.004) (0.004)
1(Turned around, 3rd) 0.009 0.009
(0.006) (0.006)
1(Turning around, 3rd) 0.049 0.049
(0.003) (0.003)
Control X X X X
IV X X
Num.Obs. 267062 267062 267062 267062
R2 0.050 0.021 0.045 0.021
R2 Adj. 0.046 0.017 0.041 0.017
RMSE 0.45 0.48 0.45 0.48
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, and federation. The standard errors in the brackets are clustered at federation categories.
Table 7 provides insights into how pressure influences the success probabilities of the second and third lifting attempts. In Column 1, the pressure from being turned around in the second attempt significantly increases the likelihood of success, as indicated by the strong positive coefficients (0.072 and 0.037). This suggests that lifters exert greater effort when facing the fear of loss aversion, possibly driven by the desire to avoid a lower rank. In contrast, Column 2 shows that the pressure of being turned around by lower-ranked lifters in the third attempt has a less pronounced effect on success, with a small positive but relatively low coefficient (0.009). This may imply that lifters exert less effort to maintain their rank in the third attempt, likely due to fatigue and the greater difficulty of improving marginal gains in ranking at this stage.
However, in both attempts, the pressure of turning around higher-ranked lifters increases the probability of success. This is evidenced by the significant positive coefficient for the third attempt (0.049), indicating that lifters are motivated to exert more effort to improve their rank when facing upward pressure. Columns 3 and 4, which present the results of the IV regression, show similar patterns. Overall, these findings highlight the shifting dynamics of pressure and motivation between the second and third attempts. Loss aversion appears to play a crucial role in the second attempt, while efforts to climb in rank are significant in both attempts, as anticipated in Table 5.
6 Counterfactual
Individual responses to pressure influence the overall distribution of risk-taking, success probability, and expected outcomes. The unique structure of official bench press competitions provides an opportunity to disentangle these effects and compare alternative competition settings to the current format.
Using the estimated results, we first simulate the attempt weights relative to the current best weight and the success probability for each attempt of each player under actual pressure. We then calculate the expected achieved weights relative to the current best, defined as the product of the attempt weight minus the current best weight and the success probability. These serve as benchmark outcomes under actual pressure conditions.
The first alternative competition setting assumes that players would have to simultaneously submit their planned attempt weights for all attempts before the first attempt and commit to the plan. This setting eliminates the possibility of pressure from rivals influencing weight attempts. While introducing this as a formal rule is unrealistic, such a commitment is common at the individual level. We refer to this scenario as the “no pressure during attempt” setting.
The second alternative competition setting posits that players could not observe the attempt weights and outcomes of their rivals after the competition begins. This removes the influence of sequentially updated outcomes from rivals on the success probability. Similar to the first scenario, formalizing this rule would be challenging, though such a commitment is feasible at the individual level. We refer to this scenario as the “no pressure during lifting” setting.
The third alternative competition setting combines the first two scenarios, creating a condition in which there would be no pressure during both the attempt and the lifting phases.
For these alternative competition settings, we simulate the attempt weights relative to the current best weight, the success probability, and the expected achieved weight for each player, holding the estimated coefficients of the pressure variables at zero.
6.1 Attempt and success probability
Panel (a) in Figure 4 illustrates the distribution of the proportional change of second and third attempt weights relative to the current best from benchmark to counterfactual “no pressure during attempt” setting. In the second attempt, the attempt weight minus the current best weight under counterfactual is significantly lower than the benchmark. For example, a lifter whose current best weight is 100kg would attempt 100kg under no pressure but 101 kg under actual pressure, as shown in the peaks around -0.2 and -0.9.111111The example shows $((100-100)-(101-100))/100=-1$. On the other hand, the third attempt distribution is more spread out, with some lifters attempting weights significantly above their benchmark. This is consistent with balancing the opposite signs of estimated coefficients of pressure. This demonstrates that lifters are more cautious in their second attempts, while third attempts tend to be riskier, with more ambitious weights.
Figure 4: Counterfactual attempt and lifting
Notes: We use estimated coefficients in Table 6 and Columns (3) and (4) in Table 7.
Panel (b) in Figure 4 displays the distribution of the proportional change of second and third attempt success probabilities from benchmark to counterfactual. The changes in success probabilities are heavily concentrated below zero for both the second and third attempts, indicating that lifters would be more likely to fail under no pressure. The second attempt distribution shows a larger concentration in the range between -0.2 and 0.02. For example, a lifter who achieves the lift by 60% in the sense of objective probability would succeed in lifting by 51 % under no pressure, as shown in the peak around -0.15.121212The example shows $(51-60)/60=-0.15$. The third attempt distribution shifts more towards zero, which means that the effect of pressure is milder in the third attempt than the second attempt. Overall, success probabilities would decrease under no pressure from rivals.
6.2 Expected weight
Figure 5: Counterfactual expected achieved weight minus current best weight
Notes: We use estimated coefficients in Table 6 and Columns (3) and (4) in Table 7.
Figure 5 presents the distribution of expected achieved weight differences from the lifters’ current best for second and third attempts, comparing the conditions of “no pressure” during the attempt, lifting, or both.
For the second attempt, the expected outcome reveals a concentration of attempts around -0.2 and -0.9, particularly in the “no pressure in both” setting, suggesting that lifters would tend to attempt weights significantly below their current best in the absence of pressure. In contrast, the “no pressure during lifting” setting displays a sharp peak near zero, indicating a large number of lifters attempting weights slightly below their personal best without pressure during the lift. These patterns suggest that, in the absence of pressure, the expected outcome for most lifters would be lower than the benchmark, largely due to less challenging attempts.
In the third attempt, the expected outcome differs markedly from the second attempt, with a greater concentration of attempts close to zero, particularly in the “no pressure in both” and “no pressure during lifting” settings. Lifters are more inclined to attempt weights near or slightly above their current best in the third attempt compared to the second. This suggests that, under no pressure, the expected outcome might exceed the benchmark for a significant portion of lifters, mainly due to more challenging attempts, albeit with a slight decrease in success probability.
As practical advice for individual lifters seeking to maximize their expected achieved weight, each lifter should consider adjusting their attempt plan based on the pressure experienced during the second attempt, while maintaining commitment to the original plan in the third attempt, disregarding pressure from rivals. This strategy is crucial for final rankings. Similarly, in broader contexts such as maximizing expected gains from risky choices, individuals should commit to their original risky decisions, ignoring external pressures from competitors.
7 Robustness check
7.1 Choice of lifter’s current best score
As our reference point, lifter’s current best, is crucial, we need to take care of the construction. For example, lifters might consider the last 12 months best score as the current best and ignore the more past best score because older personal bests may not accurately reflect lifters’ current abilities. Although we partly take care about the issues by controlling age-class, we need to check robustness of our reference point choice. We consider personal best over the more recent competition, that is, the personal best achieved within a certain recent period (e.g., the last 12 months).
Table 8: Regression of the difference between the second and third attempt weight compared to the current best within a year on pressure
(1) (2)
Dependent Variable $W_{it}^{2}-$(best) $W_{it}^{3}-$(best)
Male 0.335 7.548
(0.396) (0.734)
Body weight 0.162 0.298
(0.029) (0.040)
Num experience -0.162 -0.188
(0.077) (0.078)
1(first participation) 30.340 31.370
(1.314) (1.405)
Pressure, lower rival, 2nd 0.299
(0.007) 1
Pressure, higher rival, 2nd 0.297
Pressure, lower rival, 3rd 0.551
Pressure, higher rival, 3rd -0.100
Control X X
Num.Obs. 248778 248778
R2 0.628 0.367
R2 Adj. 0.627 0.364
RMSE 31.09 40.89
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, and federation fixed effects. The standard errors in the brackets are clustered at federation categories. $W_{it}^{2}$ and $W_{it}^{3}$ are lifter $i$’s attempt weights in the second and third attempts in competition $t$. Pressure from lower rival at 2nd is (lower rival 2nd attempt) - (his 1st outcome). Pressure from lower rival at 3rd is (lower rival 3rd attempt) - (his 2nd outcome). Pressure from higher rival at 2nd is (higher rival 2nd attempt) - (his 1st outcome). Pressure from higher rival at 3rd is (higher rival 3rd attempt) - (his 2nd outcome).
Table 9: Regression of success probability of second and third attempts on pressure, the current best within a year
(1) (2) (3) (4)
Dependent Variable 1(success 2nd) 1(success 3rd) 1(success 2nd) 1(success 3rd)
Male -0.006 0.008 -0.002 0.007
(0.006) (0.006) (0.006) (0.006)
Body weight -0.002 -0.002 -0.002 -0.002
(0.000) (0.000) (0.000) 1 (0.000)
Num experience 0.004 0.001 0.003 0.001
(0.000) (0.000) (0.000) (0.000)
1(first participation) 0.025 0.019 0.040 0.016
(0.003) (0.003) (0.003) (0.004)
$W_{it}^{2}-$(best) -0.001 -0.002
(0.000) (0.000)
$W_{it}^{3}-$(best) 0.000 0.000
(0.000) (0.000)
1(Turned around, 2nd) 0.076 0.082
(0.009) (0.009)
1(Turning around, 2nd) 0.038 0.038
(0.003) (0.004)
1(Turned around, 3rd) 0.012 0.011
(0.006) (0.006)
1(Turning around, 3rd) 0.051 0.051
(0.003) (0.003)
Control X X X X
IV X X
Num.Obs. 248778 248778 248778 248778
R2 0.047 0.020 0.044 0.020
R2 Adj. 0.042 0.016 0.040 0.016
RMSE 0.45 0.49 0.45 0.49
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, and federation. The standard errors in the brackets are clustered at federation categories.
Table 8 corresponds to Table 6 for the analysis of attempt weights using the lifter’s current best within one year. The main results remain robust under this alternative specification. Similarly, Table 9 corresponds to Table 7 for the analysis of success probabilities, utilizing the lifter’s current best within a year. Most of the main results are consistent when using this alternative measure. The only notable difference is that the coefficient for the dummy variable indicating “turning around the higher-ranked rival” shows a statistically significant positive effect.
7.2 Expected rivals’ attempt weights
In our theoretical prediction in Section 3 and empirical exercises, we treat the attempt weight of lifter $i$’s higher rival, which is not observable to lifer $i$ in his attempt stage, as the actual attempt weight in a rational expectation manner. Although this operation is ad hoc, we believe that treating the higher rival’s attempt weight as completely unknown information ignores pressure from the higher rank rival, which matters in the actual competition.
Table 10 presents the Root Mean Squared Error (RMSE) and $R^{2}$ values for various prediction models of attempt weights using linear regression. Predictive accuracy is assessed through 5-fold cross-validation. All models include gender, body weight, number of competition experiences, a dummy variable for first participation, current best, and fixed effects for equipment category, age class, division, weight class, and federation as predictors. In the full model for the 2nd attempt, both the attempt weight and the success of the 1st attempt are included as predictors, while in the “without 1st success” model, the success of the 1st attempt—which is unobservable when predicting the higher rival’s 2nd attempt weight—is excluded. Similarly, in the full model for the 3rd attempt, the attempt weight and the success of both the 1st and 2nd attempts are used, whereas in the “without 2nd success” model, the success of the 2nd attempt—unobservable when predicting the higher rival’s 3rd attempt weight—is excluded.
The prediction for the 1st attempt yields an RMSE of 24.794 and an $R^{2}$ of 0.798, indicating that accurately predicting the 1st attempt is relatively challenging. In contrast, predictions for the 2nd and 3rd attempts across all models show an $R^{2}$ exceeding 0.99, demonstrating that high predictive accuracy can be achieved even without incorporating the outcome of the preceding attempt, which is unobservable when predicting the attempt weight of a higher-ranked rival. When comparing the full models with the “without” models for both the 2nd and 3rd attempts, the full model exhibits slightly better predictive accuracy. This suggests that the outcome of an attempt reflects the lifter’s condition on the day of the competition, and lifters choose their weights accordingly, while lower-ranked rivals are unable to observe this condition. These results support our assumption that the pressure variable, derived from predictions of the higher rival’s attempt weight, is independent of unobserved factors such as the lifter’s condition and can be considered exogenous.
Table 10: Prediction of attempt weights using linear regression
Attempt Model RMSE R2
1st Full 24.795 0.798
2nd Full 4.257 0.994
2nd Without 1st success 4.858 0.993
3rd Full 3.433 0.996
3rd Without 2nd success 4.158 0.995
Notes: All models include gender, body weight, equipment category, age class, division, weight class, federation, first participation, number of competition experiences, and current best as predictors. In the full model for the 2nd attempt, both the attempt weight and the success of the 1st attempt are included as predictors, while in the “without 1st success” model, the success of the 1st attempt is excluded. Similarly, in the full model for the 3rd attempt, the attempt weight and the success of both the 1st and 2nd attempts are used, while in the “without 2nd success” model, the success of the 2nd attempt is excluded. Predictive accuracy is assessed using 5-fold cross-validation.
8 Conclusion
Using the data from bench press competitions, our study highlights the critical impact of competitive pressure on weight attempts and success probability which are analogous to lifters’ decision-making and performance. Pressure from rivals, particularly during second attempts, leads to more aggressive weight selections, driven by loss aversion and the desire to avoid a lower rank. In contrast, third attempts are influenced by fatigue and the diminishing potential for rank improvement, resulting in more conservative behavior under pressure from higher-ranked rivals. Counterfactual simulations of no-pressure scenarios reveal that lifters tend to be more conservative in their second attempts but more ambitious in their third, though success probabilities decrease without the motivating influence of competition. These results suggest that while pressure can drive lifters to perform at their best, particularly in earlier attempts, maintaining a focus on personal goals and committing to original plans may be essential for maximizing success in later stages of competition.
As a future direction, structural estimation regarding individual lifters across attempts would be interesting. Each player allocates his stamina to three attempts strategically and dynamically, which is a more prevalent feature in the popular SBD (Squat-Benchpress-Deadlift) competition. Incorporating the sequential structure across the first, second, and third attempts for all players, structural estimation of choosing weight attempt, lifting the weight as a lottery choice, and realizing the gain under their stamina constraint would be an interesting avenue.
Appendix A Additional figures and tables
Data patterns of the weight attempt and success probability other than Figures 2 and 3 in HTML in authors’ Github. Overall, we find that lifters appear to struggle more in the third attempt, especially with larger deviations from their best, indicating a significant drop in success probability compared to the second attempt.